Use Gauss'e approach to find the following sum (Do not use formulas):
3+9+15+21+...69
The sum of the sequence is _____?

Answer by Edwin McCravy(20056)

We have these 12 terms to add:

3+9+15+21+27+33+39+45+51+57+63+69

The 1st and 12th terms have sum  3+69 = 72
The 2nd and 11th terms have sum  9+63 = 72
The 3rd and 10th terms have sum 15+57 = 72
The 4th and  9th terms have sum 21+51 = 72
The 5th and  8th terms have sum 27+45 = 72
The 6th and  7th terms have sum 33+39 = 72

That's 6 times 72 or 432.

Edwin